# Agent²: An Agent-Generates-Agent Framework for Reinforcement Learning Automation

### v1

@@ Methods
To rigorously demonstrate the effectiveness of our two-stage pipeline, we quantitatively compare the baseline (default manual configuration), after Task-to-MDP Mapping (Stage 1), and after subsequent Algorithmic Optimization (Stage 2). As summarized in Figure 8, most environments and algorithms show clear improvements: Task-to-MDP Mapping improves performance in 83% of scenarios, and Algorithmic Optimization delivers additional gains in 67% of cases.
After Stage 1, In Ant-v4 with TD3, the average episode reward jumps from 3853.8 (baseline) to 5981.4, an impressive 55% improvement. In the challenging SMAC-1c3s5z scenario, the win rate increases from 0.17 to 0.21 (a 25% relative gain), and in MetaDrive with SAC, the reward rises from 178.2 to 219.6 (a 23% increase). For most environments, Task-to-MDP Mapping alone enables the agent to outperform the baseline by a large margin, demonstrating the effectiveness of automatic MDP adaptation.
Second, Algorithmic Optimization further boosts performance, especially in more challenging tasks or with sensitive algorithms. For instance, in Ant-v4 with PPO, Stage 2 brings an additional 13% improvement over Stage 1 (3831.9 vs. 3385.4). In Humanoid-v4 with SAC, performance improves by another 11.6% after Stage 2 (6787.9 vs. 6081.2). Similarly, in MetaDrive with SAC, the reward increases by an additional 18% (259.8 vs. 219.6), and in SMAC-1c3s5z, the win rate continues to climb from 0.21 to 0.23 (a further 10% relative gain).
These findings highlight the complementary nature of our pipeline: Task-to-MDP Mapping accurately constructs a specialized training MDP from natural language task descriptions, enabling rapid adaptation of the agent to the environment structure, while Algorithmic Optimization fine-tunes learning for maximum effectiveness. Together, these stages enable robust and fully automated generation of high-performing RL agents across a diverse range of single-agent and multi-agent benchmarks.
Figure 8: Performance improvement comparison across Task-to-MDP Mapping and Algorithmic Optimization.
Conclusion
In this work, we present Agent2, a fully automated, LLM-driven agent-generates-agent framework that significantly advances the automation, accessibility, and performance of RL development. By introducing a novel dual-agent architecture and a two-stage pipeline that separates MDP modeling from algorithmic optimization, Agent2 is able to autonomously translate high-level task descriptions and raw environment code into high-performing RL agents—entirely without human intervention.
Extensive experiments across single-agent and multi-agent benchmarks demonstrate that Agent2 consistently outperforms manually designed baselines, achieving performance improvements across all evaluated environments and algorithms, which validates both the effectiveness and generality of the proposed approach. Our ablation analysis further highlights the complementary impact of automated environment adaptation and iterative algorithmic refinement. Notably, Agent2 achieves performance improvement in 83% of task-algorithm pairs after automated MDP adaptation, and a further gain in 67% of cases following algorithmic optimization.
Ultimately, Agent2 establishes a new paradigm in which intelligent agents can autonomously design, optimize, and deploy other agents, thereby breaking through traditional barriers of RL development and setting the stage for the next generation of automated AI systems.

### v2

@@ Appendix B Experimental Results
To demonstrate the effectiveness of our two-stage pipeline, we quantitatively compare the baseline (default manual configuration), after Task-to-MDP Mapping (Stage 1), and after subsequent Algorithmic Optimization (Stage 2). The detailed results are summarized in Table 4.
Table 4: Performance comparison on 2 stages
Environment Scenario Algorithm Baseline Stage1 Stage2
MuJoCo Ant PPO 3280.51 3385.38 3831.86
MuJoCo Ant TD3 3853.84 5981.39 5981.39
MuJoCo Ant SAC 3499.76 3960.73 4372.05
MuJoCo Humanoid PPO 917.64 1085.85 1085.85
MuJoCo Humanoid TD3 354.79 5425.46 5425.46
MuJoCo Humanoid SAC 4682.83 6081.16 6787.94
MuJoCo Hopper PPO 3304.03 3314.88 3444.56
MuJoCo Hopper TD3 3486.29 3570.13 3570.13
MuJoCo Hopper SAC 3517.41 3517.41 3576.54
MuJoCo Walker2d PPO 4236.68 4649.51 4649.51
MuJoCo Walker2d TD3 4513.16 4844.38 4844.38
MuJoCo Walker2d SAC 4730.53 4730.53 4805.28
MetaDrive metadrive PPO 245.34 264.57 277.54
MetaDrive metadrive SAC 178.16 219.60 259.76
MPE simple_spread_v3 MAPPO -19.73 -16.63 -16.31
MPE simple_reference_v3 MAPPO -19.61 -19.00 -19.00
SC2 8m MAPPO 0.77 0.92 0.94
SC2 2s3z MAPPO 0.82 0.84 0.85
SC2 1c3s5z MAPPO 0.17 0.21 0.23